OpenStudy (anonymous):

The expression 4x-3 will equal 5 when x equals which of the following?

OpenStudy (anonymous):

4x-3=5 4x-3+3=5+3 4x=8 4x/4=8/4 x=2 Verify: 4(2)-3=8-3=5, answer is correct. 4x-3 will equal 5 when x equals 2